Thursday is ladies night at the slurp and burp bar and grill all adult beverages are $1.75 for women and $2.25 for men a total of 714 adult beverages were sold last Thursday night if the slurp and burp sold a total of $1416 in adult beverages last Thursday night how many adults beverages results in women?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52803)

.
Thursday is ladies night at the slurp and burp bar and grill all adult beverages are $1.75 for women and $2.25 for men
a total of 714 adult beverages were sold last Thursday night if the slurp and burp sold a total of $1416
in adult beverages last Thursday night how many adults beverages results in women?
~~~~~~~~~~~~~~~~~~~~

Let     x      be the number of the women drinks;
Then  the number of the   men drinks is (714-x).


Write the total money equation

    2.25*(714-x) + 1.75*x = 1416  dollars.


From this equation, get

    2.25*714 - 2.25x + 1.75x = 1416

    2.25*714 - 1416 = 2.25x - 1.75x

          190.5     =     0.5x

          x         =     190.5/0.5 = 381.


ANSWER.  381 women drinks.


CHECK.   2.25*(714-381) + 1.75*381 = 1416 dollars, total money.   ! correct !

Solved.



Answer by greenestamps(13200)



The other tutor showed a typical formal algebraic solution.

If formal algebra is not required, here is a solution using a method that can be used on any problem like this.

(1) If all 714 beverages were for men, the total cost would have been 714($2.25) = $1606.50.
(2) The actual total cost was $1416, which is $1606.50-$1416 = $190.50 less.
(3) Since each beverage for a woman costs $0.50 less than each beverage for a man, the number of beverages for women was $190.50/$0.50 = 381.

ANSWER: 381